I. Experiment V: Equilibrium Constant Lab Report ( 50 pts) Calibration Curve 1. Record the absorbance for each of the standar2. Graph Absorbance vs. [Fe(SCN)24] to generate your calibration curve for this experiment using Excel. Ensure that all axesII. Determination of the equilibrium constant 1. Record the absorbance values for each of the equilibrium solutions in par [FCan someone double check if the (Fe(SNC)2+ was calculated correctly for the B concentrations? calibration curve equation was y=2834x+0.0293. Feel free to please double check the work from A as well
Answer #1

you made mistake in A-6 . so all equations changed due to this

Ewe know that Concentratim orole volume And we also know that M = 0.002 m MV = m ₂ So for At Vio m = 0,002 m₂ = 3 ox0.002= 1LENO for A-4 = 9 m2 2 for A-s mv = M₂ V2 N = Ilmi my=Mh 0.0028 g = My x 100 0.002X11 - My XL62 Me = 1.88104 I m = 2.28104 forB-2 0.160= 1299.87X + 0.33576 X2 0-180-0.33576 1299.87 1 x = . 1-198104 0.348 = 1299.87X4 0.33576 Xz 0.348- 0.335768 1299.87